OpenStudy (anonymous):

Which expression is a sixth root of -1+i√3? a.2^(1/6)(cos(300) + isin(300)) b.2^(1/6)(cos(60) + isin(60)) c.2^(1/6)(cos(90) + isin(90)) d.2^(1/6)(cos(20) + isin(20))

OpenStudy (dumbcow):

the solution is \[2^{1/6} (\cos \frac{120 +360k}{6} + i \sin \frac{120 +360k}{6})\] for k = 0,1,2,3,4,5